OpenStudy (anonymous):

Need answer for this asapSolve for x: −3|x + 7| = −12 please.

OpenStudy (faiqraees):

square both sides

OpenStudy (anonymous):

can i just have the answer for this one, this is the only one i have left to do @FaiqRaees

OpenStudy (texaschic101):

- 3 |x + 7 | = - 12...divide both sides by - 3 | x + 7 | = 4 x + 7 = 4 - (x + 7 ) = 4 x = 4 - 7 - x - 7 = 4 x = - 3 - x = 11 x = - 11 so x = -3 and - 11